Question 5 Not yet answered Marked out of 1 P Flag question What is the total happiness of someone who plays the following decision tree with a short-sighed approach? 10 Now Answer
Question 6 Not yet answered Marked out of 1 P Flag question What is the total actual happiness of someone who plays the following decision tree with a biased and long-sighed approach? 5% 21 4 715 70% 412 60%, D B : 40% 513 Now 5 1 2
Answer #1

Answer:

Question 5)

In short sighted approach we will see only immediate benefits not long term benefits.

Like between A and B , we choose B , because it has higher value(4) compare to A(2).

After B we have to choose between D & E, we will choose E ,because it has higher value(3) compare to D(2).

After E we have to choose between H & I, we will choose H ,because it has higher value(3) compare to I(1).

So total happiness will be for B-E-H=4+3+3=10